# Tidal Love numbers and Green’s functions in black hole space-times

### v1

@@ V.2 Kramers-Kronig relations
These are the well-known Kramers-Kronig relations. They follow from causality and show that the dissipative (imaginary) part of the response function is determined in terms of the conservative (real) part, and vice versa. The relationship is, however, non-local in frequency space, as it requires knowledge of one part of the Green’s function across all frequencies to reconstruct the other for any individual frequency.
@@ VI The special case: BTZ black holes
The purpose of this Section is to study a simple case where TLNs and the Green’s function can be directly related in general relativity; namely the class of BTZ black hole solutions in 2+1 dimensional gravity Banados et al. (1992, 1993). These solutions are special since gravity in $2+1$ dimensions is non-dynamical, hence there are no radiative modes.
Non-extremal rotating BTZ BHs are solutions to Einstein’s equations with a negative cosmological constant $\Lambda=-1/l^{2}$. Their metric, in $(t,r,\varphi)$ coordinates, reads
@@
where the the inner and outer BH radii are given by
@@
Consider a minimally-coupled massless scalar field propagating on this background. Adopting the standard decomposition
@@
To draw another contrast, the behavior of the BTZ effective potential (71) at spatial infinity does not induce a late-time power law tail in the Green’s function. This is unlike the power-law tail in asymptotically flat space-times, as seen in Eq. (59). Instead, the retarded Green’s function for BTZ BHs decays exponentially Chan and Mann (1997); Horowitz and Hubeny (2000). This is compatible with the hypergeometric solutions to the wave equation, which do not show any branch cut in the complex frequency plane (see related discussion in Sec. V). The absence of tail effects and gravitational waves allowed us to derive an explicit, closed-form relation between TLNs and retarded Green’s function.
@@ VII Conclusions
The purpose of this paper is two-fold. Firstly, we have highlighted the intrinsic connection between dynamical response and the retarded Green’s function. Secondly, we have demonstrated the ambiguity in extracting the former from the latter when radiation reaction and related phenomena are present.
For asymptotically flat BHs, the retarded Green’s function receives three contributions: $i)$ from high-frequency modes propagating directly from the source to the observer; $ii)$ from quasinormal modes emitted during the relaxation to equilibrium of the perturbed BHs; and $iii)$ from tail effects induced by the backscattering of the emitted gravitational waves on the asymptotic curved space-time. As we have seen, the presence of tail effects, which provide an example of radiation reaction on the system induced by physics at large distances, contaminates the extraction of dynamical Love numbers.
@@
This work represents an initial step towards a full investigation of the relation between tidal interactions and the retarded Green’s function. It can be improved in various directions. Firstly, it would be important to investigate alternative ways of solving the wave equation within the buffer region, based on analytical or numerical approaches, to have a more in-depth investigation of dynamical TLNs and dissipative coefficients. Secondly, it would be interesting to perform a more detailed study to understand how radiation reaction and tail effects enter the solution to the wave equation, in order to possibly disentangle their contribution from the tidal one. Finally, the complete knowledge of the retarded Green’s function could also allow us to explicitly check the Kramers-Kronig relations for BHs, which could provide additional insights into the interplay between the near and wave zone physics. We leave these studies for future work.

### v2

@@ V.2 Kramers-Kronig relations
These are the well-known Kramers-Kronig relations. They follow from causality and show that the dissipative (imaginary) part of the response function is determined in terms of the conservative (real) part, and vice versa. The relationship is, however, nonlocal in frequency space, as it requires knowledge of one part of the Green’s function across all frequencies to reconstruct the other for any individual frequency.
@@ VI The special case: BTZ black holes
The purpose of this section is to study a simple case where TLNs and the Green’s function can be directly related in general relativity; namely the class of BTZ black hole solutions in 2+1 dimensional gravity Banados et al. (1992, 1993). These solutions are special since gravity in $2+1$ dimensions is nondynamical, hence there are no radiative modes.
Nonextremal rotating BTZ BHs are solutions to Einstein’s equations with a negative cosmological constant $\Lambda=-1/l^{2}$. Their metric, in $(t,r,\varphi)$ coordinates, reads
@@
where the inner and outer BH radii are given by
@@
Consider a minimally coupled massless scalar field propagating on this background. Adopting the standard decomposition
@@
To draw another contrast, the behavior of the BTZ effective potential (71) at spatial infinity does not induce a late-time power law tail in the Green’s function. This is unlike the power-law tail in asymptotically flat spacetimes, as seen in Eq. (59). Instead, the retarded Green’s function for BTZ BHs decays exponentially Chan and Mann (1997); Horowitz and Hubeny (2000). This is compatible with the hypergeometric solutions to the wave equation, which do not show any branch cut in the complex frequency plane (see related discussion in Sec. V). The absence of tail effects and gravitational waves allowed us to derive an explicit, closed-form relation between TLNs and retarded Green’s function.
@@ VII Conclusions
The purpose of this paper is two-fold. First, we have highlighted the intrinsic connection between dynamical response and the retarded Green’s function. Second, we have demonstrated the ambiguity in extracting the former from the latter when radiation reaction and related phenomena are present.
For asymptotically flat BHs, the retarded Green’s function receives three contributions: $i)$ from high-frequency modes propagating directly from the source to the observer; $ii)$ from quasinormal modes emitted during the relaxation to equilibrium of the perturbed BHs; and $iii)$ from tail effects induced by the backscattering of the emitted gravitational waves on the asymptotic curved spacetime. As we have seen, the presence of tail effects, which provide an example of radiation reaction on the system induced by physics at large distances, contaminates the extraction of dynamical Love numbers.
@@
This work represents an initial step toward a full investigation of the relation between tidal interactions and the retarded Green’s function. It can be improved in various directions. First, it would be important to investigate alternative ways of solving the wave equation within the buffer region, based on analytical or numerical approaches, to have a more in-depth investigation of dynamical TLNs and dissipative coefficients. A promising strategy would be to consider higher-order solutions in the MST method, which has proven to be a powerful approach for solving the Teukolsky equation. Second, it would be interesting to perform a more detailed study to understand how radiation reaction and tail effects enter the solution to the wave equation, in order to possibly disentangle their contribution from the tidal one. Finally, the complete knowledge of the retarded Green’s function could also allow us to explicitly check the Kramers-Kronig relations for BHs, which could provide additional insights into the interplay between the near and wave zone physics. We leave these studies for future work.